OpenStudy (anonymous):

given linear equation y=-3/2x+1 find the y- coordinates of the points(-4), (-2) and (6). Please show all of your work. Plot those points and graph the linear equation.

OpenStudy (anonymous):

For x = -4: y=-3/2(-4)+1 y=12/2+1 y=6+1 y=7 For x = -2: y=-3/2(-2)+1 y=6/2+1 y=3+1 y=4 For x = 6: y=-3/2(6)+1 y=-18/2+1 y=-9+1 y=-8 From the equation, it can be seen the y-int is 1. From the equation, it can be seen the x-int is 2/3.